# Beyond Nash-Williams: Counterexamples to Clique Decomposition Thresholds for All Cliques Larger than Triangles

### v1

@@ 1. Introduction
Nash-Williams’ Conjecture may be seen as the foundational problem in what may be termed extremal design theory; it is instructive to compare this value to the fractions in the minimum degree thresholds for other extremal problems: namely the threshold is $1/2$ for the existence of at least one triangle (which follows from a classical theorem usually attributed to Mantel [25] from 1907) and $2/3$ for the existence of a triangle factor, that is a set of vertex-disjoint triangles spanning all vertices (as given by the celebrated Corrádi-Hajnal Theorem [9]). The current best-known results on Nash-Williams’ Conjecture are as follows: In 2016, Barber, Kühn, Lo, and Osthus [2] proved that if the fractional relaxation of Nash-Williams’ Conjecture holds for minimum degree $cn$ for some constant $c\geq 3/4$, then Nash-Williams’ Conjecture holds for any constant $c^{\prime}>c$; combined with the best-known fractional result of Delcourt and Postle [12] from 2021, this confirms that Nash-Williams’ Conjecture holds for graphs with minimum degree at least $0.82733n$. This improved a series of earlier works on the fractional relaxation by Garaschuk [18] in 2014, Dross [13] in 2016, and Dukes and Horsley [16] in 2020. Nash-Williams’ Conjecture has become a central open question in design theory as various generalizations are now tied to it (e.g., the minimum degree threshold for an $F$-decomposition for any $3$-chromatic $F$ [19], the minimum degree threshold for approximately packing families of $2$-regular graphs [8], and the minimum degree threshold for high-girth triangle decompositions [11]). For more history on Nash-Williams’ Conjecture and its generalizations, see the survey by Glock, Kühn and Osthus [20].
@@ 1.1. Decompositions of $q$-partite graphs
We should note that Montgomery [27] specifically singled out the fractional relaxation of Conjecture 1.6 as interesting in its own right. That said, Montgomery [27] noted that the fractional relaxation of Conjecture 1.6 implies the fractional relaxation of Conjecture 1.2, and indeed furthermore if the former held for partite minimum degree $cn$, then the later would hold for minimum degree $cn$. Thus it follows from Theorem 1.3 that we immediately disprove Conjecture 1.6 for all $q\geq 4$ as follows. We explain below the connection between these two conjectures for the interested reader.
@@
Here is the connection observed by Montgomery [27]. Recall that the tensor product of two graphs $G_{1}$ and $G_{2}$, denoted $G_{1}\times G_{2}$, is the graph with $V(G_{1}\times G_{2})=V(G_{1})\times V(G_{2})$ where two vertices $(v_{1},v_{2})$ and $(u_{1},u_{2})$ are adjacent if and only if $v_{1}u_{1}\in E(G_{1})$ and $v_{2}u_{2}\in E(G_{2})$. Notice that the graph $G\times K_{q}$ is a balanced $q$-partite graph with parts of size $v(G)$ and that $\hat{\delta}(G\times K_{q})=\delta(G)$. Moreover $G\times K_{q}$ is partite-$K_{q}$-divisible for any $G$ by construction. Finally, we note that $G\times K_{q}$ admits a fractional $K_{q}$-decomposition if and only if $G$ does. (The backward direction follows by using the same weight for a copy of $K_{q}$ in $G\times K_{q}$ as the weight of its original copy in $G$ while the forward direction follows by using as the weight of a clique in $G$ the average of all its partite copies in $G\times K_{q}$.)
@@ 2.1. Blow-ups: Reducing to one counterexample for each $q$
Outcome (a) is immediate from the construction. Outcome (b) follows since the degree of each vertex of $G\circ\overline{K_{m}}$ is divisible by $m$, and hence by q1q-1, and in addition we also then find that $e(G\circ\overline{K_{m}})$ is divisible by $\frac{m}{2}$, and hence by $\binom{q}{2}$. For outcome (c), the backward direction follows by using the same weight for a copy of $K_{q}$ in $G\circ\overline{K_{m}}$ as the weight of its original copy in $G$ while the forward direction follows by using for the weight of a clique in $G$ the average of all its corresponding copies in $G\circ\overline{K_{m}}$. ∎
@@ 2.2. Intuition for the constructions
It may be useful to the reader to understand that there have been, in fact, two different constructions in the literature showing that the value of $1-\frac{1}{q+1}$ in Conjecture 1.2 would have been tight if true. The first construction due, to Gustavsson in 1991 [21] on pages 85-86 of his thesis (and reiterated by Yuster [36]), is $K_{q-1}\circ H$ where $H$ is a $d$-regular graph on $v$ vertices for some specific $d$ and $v$ with the property that $d/v$ is close to but smaller than $2/(q+1)$. The second construction, due to Garaschuk in 2014 [18] (and reiterated by Barber, Kühn, Lo, and Osthus [2]), is $(K_{q+1}-M)\circ K_{t}$ (the $t$-clique-blowup) where $M$ is a perfect matching of $K_{q+1}$ (for $q$ odd). In both constructions, one can show that no fractional $K_{q}$-decomposition exists by arguing there are not enough ‘internal’ edges to decompose the ‘cross’ edges.
@@
In our general construction we fix some $d$-regular graph $H$ on $v$ vertices and then consider $G:=K_{b}\circ H$, where $b$ is some integer to be specified (equivalently $G$ is the join of $b$ copies of $H$). We note the particular graph $H$ is not important to the construction but rather ultimately only the parameters $d$, $v$, and $b$ matter. We call an edge of $K_{b}\circ H$ between distinct copies of $H$ a cross edge and an edge of $K_{b}\circ H$ that lies inside a copy of $H$ an internal edge. Note there are exactly $b\cdot\frac{dv}{2}$ internal edges of $K_{b}\circ H$ and exactly $\binom{b}{2}\cdot v^{2}$ cross edges of $K_{b}\circ H$. The important observation is the following: if $b<q$, then every copy of $K_{q}$ in $K_{b}\circ H$ contains at least qbq-b internal edges. This observation follows from the pigeonhole principle (indeed by the convexity of binomial coefficients it follows that the minimum of internal edges is attained when the vertices of a copy of $K_{q}$ are distributed as equitably as possible. This observation implies that internal edges can only decompose (even fractionally) cross edges at a ratio of (qb)(q-b) to $\binom{q}{2}-(q-b)$. However, if the ratio of internal edges to cross edges, which is equal to $\frac{(bdv)/2}{\binom{b}{2}v^{2}}=\frac{d}{v}\cdot\frac{1}{b-1}$ is smaller than this ratio, then we see that $K_{b}\circ H$ does not admit a fractional $K_{q}$-decomposition, i.e., when
@@
The inquisitive reader may wonder what happens with our construction for $b<\frac{q}{2}$; these values of $b$ still provide counterexamples to Conjecture 1.2 but not as large a minimum degree value as obtained when $b\approx\frac{q}{\sqrt{2}}$. Suppose that $b=\frac{q}{k+\alpha}$ where $k\geq 1$ is integer and $\alpha\in[0,1]$. A stronger pigeonholing argument shows that much more than qbq-b internal edges are used for $k\geq 2$; as mentioned before, the minimum value of internal edges is achieved when the vertices of a copy of $K_{q}$ are distributed as equitably as possible among the parts, that is, when using $k+1$ vertices from $\alpha\cdot b$ parts and $k$ vertices from $(1-\alpha)\cdot b$ parts. Hence at least $\alpha\cdot b\cdot\binom{k+1}{2}+(1-\alpha)\cdot b\cdot\binom{k}{2}=(k+2\alpha-1)\cdot\frac{kb}{2}$ internal edges are used. Call this number $\gamma$. It follows that $K_{b}\circ H$ admits no fractional $K_{q}$-decomposition if $\frac{d}{v}<(b-1)\cdot\frac{\gamma}{\binom{q}{2}-\gamma}$. For large $q$, this is approximately equal to $b\cdot\frac{\gamma}{q^{2}/2}=(b^{2}/2)\cdot\frac{(k+2\alpha-1)\cdot k}{q^{2}/2}=\frac{(k+2\alpha-1)\cdot k}{(k+\alpha)^{2}}$. This in turn yields a minimum degree fraction of $1-\frac{1-\frac{(k+2\alpha-1)\cdot k}{(k+\alpha)^{2}}}{b}=1-\frac{(k+\alpha)^{2}-(k+2\alpha-1)\cdot k}{(k+\alpha)\cdot q}=1-\frac{k+\alpha^{2}}{(k+\alpha)\cdot q}$. To maximize this, we minimize $\frac{k+\alpha^{2}}{k+\alpha}=1-\frac{\alpha\cdot(1-\alpha)}{k+\alpha}$ for $\alpha\in[0,1]$ and $k$ fixed. Clearly setting $k$ smaller for any fixed $\alpha$ will only make this value smaller (so the $k=1$ ‘interval’ dominates all others); meanwhile for $k\geq 2$, the right term is at most $\frac{1/4}{2}\leq\frac{1}{8}$ leading to a value no better than $\frac{8}{7}\cdot q\approx 1.14\cdot q$ in the denominator (which is smaller than the approximately $1.207q$ in Theorem 2.3). Nevertheless for any integer $k\geq 1$ and any real $\alpha\in(0,1)$, the value $\frac{k+\alpha^{2}}{k+\alpha}$ is strictly less than $1$ and so would supply counterexamples to Conjecture 1.2 for $q$ large enough (even multiplicatively), see Figure˜1 for the constant $c=c(k,\alpha)$ obtained. Viewed holistically from this asymptotic view then, all the “integral” ratios for number of parts (the harmonies) yield a minimum degree value of $1-\frac{1}{q}$ asymptotically while the “non-integral” ratios (the discordances) yield asymptotically larger minimum degree values with the largest such value lying in the first harmonic interval at a value of $\beta\approx\frac{1}{\sqrt{2}}$ (probably to the great dismay of the Pythagorean school).
@@ 3. The General Construction and Large $q$
Here is the lemma codifying when our general construction does not admit a fractional $K_{q}$-decomposition.
@@
Let $H$ be a $d$-regular bipartite graph on $v$ vertices, and let $G:=K_{b}\circ H$. Recall that we say that an edge inside a copy of $H$ is an internal edge , while all other edges are cross edges. Observe that $G$ contains $\frac{b\cdot d\cdot v}{2}$ internal edges, and $\binom{b}{2}v^{2}$ cross edges. Assume that there exists a fractional $K_{q}$-decomposition $\phi$ of $G$. Recall that we observed earlier by the pigeonhole principle that every copy of $K_{q}$ in $K_{b}\circ H$ contains at least qbq-b internal edges. Using this, we obtain by counting for the internal edges of $G$ that
@@
$\frac{b\cdot d\cdot v}{2}\cdot\Big{[}\binom{q}{2}-(q-b)\Big{]}\cdot\geq\binom{b}{2}v^{2}\cdot(q-b).$
@@ 2 $\iff q^{3}-3\cdot b\cdot q^{2}+q\cdot(2\cdot b^{2}+b-1)<0$
The Left-Hand Side of this inequality is a (convex) quadratic polynomial whose unique root is b=q1b=q-1 when $q=3$, but has two distinct roots for $q\geq 4$, namely $b_{+}:=q-1$ and $b_{-}:=\frac{q+1}{2}$. Observe that $|b_{+}-b_{-}|=1$ for $q=5$, and $|b_{+}-b_{-}|>1$ for $q\geq 6$. Therefore for any $q\geq 6$, there exists an integer $b$ (e.g. b=q2b=q-2) satisfying ˜2, as desired. ∎
@@ 3.2. Large estimates
hence $\delta(G)\geq\left(1-\frac{1}{\Big{(}\frac{1+\sqrt{2}}{2}-\varepsilon\Big{)}\cdot(q+1)}\right)n$, as desired, ∎
@@ 4.1. The case $q=5$
Observe that $G$ contains all cross edges between copies of $H$, that is, $3v^{2}$ cross edges; each copy of $H$ is $(d\cdot v)$-regular on $v$ vertices, therefore $G$ contains $\frac{3\cdot d\cdot v^{2}}{2}$ internal edges; finally $G$ contains all small edges between $I$ and every copy of $H$, that is, $3\alpha v^{2}$ small edges. Since $I$ is independent, we have that for any $F\in\binom{G}{K_{q}}$, $|V(F)\cap V(I)|\leq 1$. For $i\in\{0,1\}$, let $\mathcal{F}_{i}$ be the family of all copies of $K_{q}$ in $G$ intersecting $I$ on exactly $i$ vertices, that is,
@@
For every $F\in\mathcal{F}_{0}$, we note that $F$ contains at least $2$ internal edges and at most $8$ cross edges (and no small edges); on the other hand, for every $F\in\mathcal{F}_{1}$, we note that $F$ contains at least $1$ internal edge, at most $5$ cross edges, and exactly $4$ small edges. As $\phi$ is a fractional $K_{q}$-decomposition of $G$, we have that $\partial\phi(e)=1$ for every edge $e\in G$. In particular we have
@@
Recall that $G$ contains $\frac{3\cdot d\cdot v^{2}}{2}$ internal edges, therefore, with $d=(1-\alpha)$, we obtain
@@
Then, recall that every $F\in\mathcal{F}_{0}$ contains at least $2$ internal edges, while every $F\in\mathcal{F}_{1}$ contains at least $1$ internal edge, therefore

### v2

@@ Abstract.
This work was completed while the third author was at the University of Waterloo.
@@ 1. Introduction
Nash-Williams’ Conjecture may be seen as the foundational problem in what may be termed extremal design theory; it is instructive to compare this value to the fractions in the minimum degree thresholds for other extremal problems: namely the threshold is $1/2$ for the existence of at least one triangle (which follows from a classical theorem usually attributed to Mantel [25] from 1907) and $2/3$ for the existence of a triangle factor, that is a set of vertex-disjoint triangles spanning all vertices (as given by the celebrated Corrádi-Hajnal Theorem [9]). The current best-known results on Nash-Williams’ Conjecture are as follows: In 2016, Barber, Kühn, Lo, and Osthus [2] proved that if the fractional relaxation of Nash-Williams’ Conjecture holds for minimum degree $cn$ for some constant $c\geq 3/4$, then Nash-Williams’ Conjecture holds for any constant $c^{\prime}>c$; combined with the best-known fractional result of Delcourt and Postle [12] from 2021, this confirms that Nash-Williams’ Conjecture holds for graphs with minimum degree at least $0.82733n$. This improved a series of earlier works on the fractional relaxation by Garaschuk [18] in 2014, Dross [13] in 2016, and Dukes and Horsley [16] in 2020. Nash-Williams’ Conjecture has become a central open question in design theory as various generalizations are now tied to it (e.g., the minimum degree threshold for an $F$-decomposition for any $3$-chromatic $F$ [19], the minimum degree threshold for approximately packing families of $2$-regular graphs [8], and the minimum degree threshold for high-girth triangle decompositions [11]). For more history on Nash-Williams’ Conjecture and its generalizations, see the survey by Glock, Kühn, and Osthus [20].
@@ 1.1. Decompositions of $q$-partite graphs
We should note that Montgomery [26] specifically singled out the fractional relaxation of Conjecture 1.6 as interesting in its own right. That said, Montgomery [26] noted that the fractional relaxation of Conjecture 1.6 implies the fractional relaxation of Conjecture 1.2, and indeed furthermore if the former held for partite minimum degree $cn$, then the latter would hold for minimum degree $cn$. Thus it follows from Theorem 1.3 that we immediately disprove Conjecture 1.6 for all $q\geq 4$ as follows. We explain below the connection between these two conjectures for the interested reader.
@@
Here is the connection observed by Montgomery [26]. Recall that the tensor product of two graphs $G_{1}$ and $G_{2}$, denoted $G_{1}\times G_{2}$, is the graph with $V(G_{1}\times G_{2})=V(G_{1})\times V(G_{2})$ where two vertices $(v_{1},v_{2})$ and $(u_{1},u_{2})$ are adjacent if and only if $v_{1}u_{1}\in E(G_{1})$ and $v_{2}u_{2}\in E(G_{2})$. Notice that the graph $G\times K_{q}$ is a balanced $q$-partite graph with parts of size $v(G)$ and that $\hat{\delta}(G\times K_{q})=\delta(G)$. Moreover $G\times K_{q}$ is partite-$K_{q}$-divisible for any $G$ by construction. Finally, we note that $G\times K_{q}$ admits a fractional $K_{q}$-decomposition if and only if $G$ does. (The backward direction follows by using for a copy of $K_{q}$ in $G\times K_{q}$ the weight of its original copy in $G$ scaled down by a factor of (q2)!(q-2)!, while the forward direction follows by using as the weight of a clique in $G$ the average of all its partite copies in $G\times K_{q}$.)
@@ 2.1. Blow-ups: Reducing to one counterexample for each $q$
Outcome (a) is immediate from the construction. Outcome (b) follows since the degree of each vertex of $G\circ\overline{K_{m}}$ is divisible by $m$, and hence by q1q-1, and in addition we also then find that $e(G\circ\overline{K_{m}})$ is divisible by $\frac{m}{2}$, and hence by $\binom{q}{2}$. For outcome (c), the backward direction follows by using for a copy of $K_{q}$ in $G\circ\overline{K_{m}}$ the weight of its original copy in $G$ scaled down by a factor of $m^{q-2}$, while the forward direction follows by using for the weight of a clique in $G$ the average of all its corresponding copies in $G\circ\overline{K_{m}}$. ∎
@@ 2.2. Intuition for the constructions
It may be useful to the reader to understand that there have been, in fact, two different constructions in the literature showing that the value of $1-\frac{1}{q+1}$ in Conjecture 1.2 would have been tight if true. The first construction due to Gustavsson in 1991 [21] on pages 85-86 of his thesis (and reiterated by Yuster [36]), is $K_{q-1}\circ H$ where $H$ is a $d$-regular graph on $v$ vertices for some specific $d$ and $v$ with the property that $d/v$ is close to but smaller than $2/(q+1)$. The second construction, due to Garaschuk in 2014 [18] (and reiterated by Barber, Kühn, Lo, and Osthus [2]), is $(K_{q+1}-M)\circ K_{t}$ (the $t$-clique-blowup) where $M$ is a perfect matching of $K_{q+1}$ (for $q$ odd). In both constructions, one can show that no fractional $K_{q}$-decomposition exists by arguing there are not enough ‘internal’ edges to decompose the ‘cross’ edges.
@@
In our general construction we fix some $d$-regular graph $H$ on $v$ vertices and then consider $G:=K_{b}\circ H$, where $b$ is some integer to be specified (equivalently $G$ is the join of $b$ copies of $H$). We note the particular graph $H$ is not important to the construction but rather ultimately only the parameters $d$, $v$, and $b$ matter. We call an edge of $K_{b}\circ H$ between distinct copies of $H$ a cross edge and an edge of $K_{b}\circ H$ that lies inside a copy of $H$ an internal edge. Note there are exactly $b\cdot\frac{dv}{2}$ internal edges of $K_{b}\circ H$ and exactly $\binom{b}{2}\cdot v^{2}$ cross edges of $K_{b}\circ H$. The important observation is the following:
if $b<q$, then every copy of $K_{q}$ in $K_{b}\circ H$ contains at least qbq-b internal edges.
This observation follows from the pigeonhole principle (indeed by the convexity of binomial coefficients it follows that the minimum of internal edges is attained when the vertices of a copy of $K_{q}$ are distributed as equitably as possible. This observation implies that internal edges can only decompose (even fractionally) cross edges at a ratio of (qb)(q-b) to $\binom{q}{2}-(q-b)$. However, if the ratio of internal edges to cross edges, which is equal to $\frac{(bdv)/2}{\binom{b}{2}v^{2}}=\frac{d}{v}\cdot\frac{1}{b-1}$ is smaller than this ratio, then we see that $K_{b}\circ H$ does not admit a fractional $K_{q}$-decomposition, i.e., when
@@
The inquisitive reader may wonder what happens with our construction for $b<\frac{q}{2}$; these values of $b$ still provide counterexamples to Conjecture 1.2 but not as large a minimum degree value as obtained when $b\approx\frac{q}{\sqrt{2}}$. Suppose that $b=\frac{q}{k+\alpha}$ where $k\geq 1$ is integer and $\alpha\in[0,1]$. A stronger pigeonholing argument shows that much more than qbq-b internal edges are used for $k\geq 2$; as mentioned before, the minimum value of internal edges is achieved when the vertices of a copy of $K_{q}$ are distributed as equitably as possible among the parts, that is, when using $k+1$ vertices from $\alpha\cdot b$ parts and $k$ vertices from $(1-\alpha)\cdot b$ parts. Hence at least $\alpha\cdot b\cdot\binom{k+1}{2}+(1-\alpha)\cdot b\cdot\binom{k}{2}=(k+2\alpha-1)\cdot\frac{kb}{2}$ internal edges are used. Call this number $\gamma$. It follows that $K_{b}\circ H$ admits no fractional $K_{q}$-decomposition if $\frac{d}{v}<(b-1)\cdot\frac{\gamma}{\binom{q}{2}-\gamma}$. For large $q$, this is approximately equal to
$b\cdot\frac{\gamma}{q^{2}/2}=(b^{2}/2)\cdot\frac{(k+2\alpha-1)\cdot k}{q^{2}/2}=\frac{(k+2\alpha-1)\cdot k}{(k+\alpha)^{2}}.$
This in turn yields a minimum degree fraction of
$1-\frac{1-\frac{(k+2\alpha-1)\cdot k}{(k+\alpha)^{2}}}{b}=1-\frac{(k+\alpha)^{2}-(k+2\alpha-1)\cdot k}{(k+\alpha)\cdot q}=1-\frac{k+\alpha^{2}}{(k+\alpha)\cdot q}.$
To maximize this, we minimize $\frac{k+\alpha^{2}}{k+\alpha}=1-\frac{\alpha\cdot(1-\alpha)}{k+\alpha}$ for $\alpha\in[0,1]$ and $k$ fixed. Clearly setting $k$ smaller for any fixed $\alpha$ will only make this value smaller (so the $k=1$ ‘interval’ dominates all others); meanwhile for $k\geq 2$, the right term is at most $\frac{1/4}{2}\leq\frac{1}{8}$ leading to a value no better than $\frac{8}{7}\cdot q\approx 1.14\cdot q$ in the denominator (which is smaller than the approximately $1.207q$ in Theorem 2.3). Nevertheless for any integer $k\geq 1$ and any real $\alpha\in(0,1)$, the value $\frac{k+\alpha^{2}}{k+\alpha}$ is strictly less than $1$ and so would supply counterexamples to Conjecture 1.2 for $q$ large enough (even multiplicatively), see Figure˜1 for the constant $c=c(k,\alpha)$ obtained. Viewed holistically from this asymptotic view then, all the “integral” ratios for number of parts (the harmonies) yield a minimum degree value of $1-\frac{1}{q}$ asymptotically while the “non-integral” ratios (the discordances) yield asymptotically larger minimum degree values with the largest such value lying in the first harmonic interval at a value of $\beta\approx\frac{1}{\sqrt{2}}$ (probably to the great dismay of the Pythagorean school).
@@ 3. The General Construction and Large $q$
The following lemma codifies when our general construction does not admit a fractional $K_{q}$-decomposition. First we need some additional notation as follows. For graphs $F,G$, we denote the set of copies of $F$ in $G$ by $\binom{G}{F}$. Given an assignment of weights $\phi$ to all all copies of $F$ in $G$, $\phi:\binom{G}{F}\to\mathbb{R}$, and given an edge $e$ of $G$, we let $\partial\phi(e)$ denote the total weight of $\phi$ over the edge $e$, that is, $\partial\phi(e):=\sum_{F:e\subseteq F}\phi(F)$.
@@
Let $H$ be a $d$-regular bipartite graph on $v$ vertices, and let $G:=K_{b}\circ H$. Recall that we say that an edge inside a copy of $H$ is an internal edge , while all other edges are cross edges. Observe that $G$ contains $\frac{b\cdot d\cdot v}{2}$ internal edges, and $\binom{b}{2}v^{2}$ cross edges. Assume that there exists a fractional $K_{q}$-decomposition $\phi$ of $G$. By the pigeonhole principle, every copy of $K_{q}$ in $K_{b}\circ H$ contains at least qbq-b internal edges. We obtain by counting the internal edges of $G$ that
@@
$\frac{b\cdot d\cdot v}{2}\cdot\Big[\binom{q}{2}-(q-b)\Big]\geq\binom{b}{2}v^{2}\cdot(q-b).$
@@ 2 $\iff q^{3}-3\cdot b\cdot q^{2}+q\cdot(2\cdot b^{2}+b-1)<0$
Remark 3.3.
The left-hand side of this inequality is a (convex) quadratic polynomial whose unique root is b=q1b=q-1 when $q=3$, but has two distinct roots for $q\geq 4$, namely $b_{+}:=q-1$ and $b_{-}:=\frac{q+1}{2}$. Observe that $|b_{+}-b_{-}|=1$ for $q=5$, and $|b_{+}-b_{-}|>1$ for $q\geq 6$. Therefore as discussed in the proof overview, for any $q\geq 6$, there exists an integer $b$ (e.g. b=q2b=q-2) satisfying ˜2, as desired.
@@ 3.2. Large estimates
hence $\delta(G)\geq\left(1-\frac{1}{\Big(\frac{1+\sqrt{2}}{2}-\varepsilon\Big)\cdot(q+1)}\right)n$, as desired. ∎
@@ 4.1. The case $q=5$
Observe that $G$ contains all cross edges between copies of $H$, that is, $3v^{2}$ cross edges; each copy of $H$ is $(dv)$-regular on $v$ vertices, therefore $G$ contains $\frac{3dv^{2}}{2}$ internal edges; finally $G$ contains all small edges between $I$ and every copy of $H$, that is, $3\alpha v^{2}$ small edges. Since $I$ is independent, we have that for any $F\in\binom{G}{K_{q}}$, $|V(F)\cap V(I)|\leq 1$. For $i\in\{0,1\}$, let $\mathcal{F}_{i}$ be the family of all copies of $K_{q}$ in $G$ intersecting $I$ on exactly $i$ vertices, that is,
@@
For every $F\in\mathcal{F}_{0}$, as $H$ is bipartite and therefore triangle-free, we note that $F$ contains $2$ internal edges and $8$ cross edges (and no small edges); on the other hand, for every $F\in\mathcal{F}_{1}$, we note that $F$ contains at least $1$ internal edge, at most $5$ cross edges, and exactly $4$ small edges. As $\phi$ is a fractional $K_{q}$-decomposition of $G$, we have that $\partial\phi(e)=1$ for every edge $e\in G$. In particular we have
@@
Recall that $G$ contains $\frac{3dv^{2}}{2}$ internal edges, therefore, with $d=(1-\alpha)$, we obtain
@@
Then, recall that every $F\in\mathcal{F}_{0}$ contains $2$ internal edges, while every $F\in\mathcal{F}_{1}$ contains at least $1$ internal edge, therefore